OpenStudy (anonymous):

The results of a survey of 200 people incoming students show the following results: 75 students own a Playstation 65 students own a Xbox 55 Students own a Nintendo 25 students own at least a PlayStation and an Xbox 20 Students own at least Nintendo and a Playstation 15 students own at least an Xbox and a Nintendo 5 own all three Playstation, Xbox, Nintendo How many people don't own any video game system?

OpenStudy (anonymous):

Can someone help me out?

OpenStudy (anonymous):

hey man, lets see what you got today

OpenStudy (anonymous):

alright

OpenStudy (anonymous):

"How many people don't own any video game system?" is the question. Is this a algebra question?

OpenStudy (anonymous):

This is actually have to do with sets

OpenStudy (anonymous):

Do you understand how to set it up

OpenStudy (anonymous):

in the first 3 lines, it gives us all the information we need. P stands for playstation X stands for x box N stands for nintendo ? stands for the people that don't have any video game system 200 = 75P + 65X + 55N + ? 25 = P + x 20 = N + X 5 = P + X + N here is what I could get set up

OpenStudy (anonymous):

How do you solve it

OpenStudy (anonymous):

have you learned about systems of equations?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

great. ...mmm this one is still got me thinking, hold on..

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i'm stumped..

OpenStudy (anonymous):

Can someone help me solve this

OpenStudy (anonymous):

I think I get it now

OpenStudy (anonymous):

lets just look at nintendo first.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

wait wait sorry, ya i'm no good at this :( sorry bigtim

OpenStudy (anonymous):

I got a number of 65, does that sound right?

OpenStudy (anonymous):

not sure

OpenStudy (anonymous):

here is what I did, I reversed the order, and started subtacting one group from the next group. we are given number of kids, but what we need to know is how many total systems there are, not how many kids there are. 5 own all three Playstation, Xbox, Nintendo, 5pxn 15 students own at least an Xbox and a Nintendo, 15xn 20 Students own at least Nintendo and a Playstation, 20np 25 students own at least a PlayStation and an Xbox, 25px 55 Students own a Nintendo, 55n 65 students own a Xbox, 65x 75 students own a Playstation, 75p

OpenStudy (anonymous):

if 5 kids own all three, then I subtract that from the next three sentances. 5pxn-5= 0pxn 15xn-5 = 10xn 20np-5 = 15np 25px-5 = 20px 55n 65x 75p

OpenStudy (anonymous):

so my new set is 0pxn 10xn 15np 20px 55n 65x 75p now i subtract the new set from the the next set. 0pxn 10xn 15np 20px 55n-10n-15n = 30n 65x-10x-20x=35x 75p-20p-15p=40p so my new set now is 0pxn 0xn 0np 0px 30n 35x 40p so there are a total of 30nintendos, 35xboxes and 40 playstations.

OpenStudy (anonymous):

30 + 35 + 40 = 135 200-135 = 65 65 students don't have any.